Can you help with this? I was told I could substitute the given numbers to get the equation, but what equation would I use?
Find the equation of the parabola (with horizontal axis of symmetry) containing the point (20,4) with vertex (4,2)
Thanks!

Answer by scott8148(6628)

horizontal axis of symmetry means a basic parabola of x=y^2 which has a vertex at the origin

the offset of the vertex results in x=a(y-2)^2+4

using the point to find a gives 20=a(4-2)^2+4 ... 16=4a ... 4=a

so the equation is x=4(y-2)^2+4 ... or x=4y^2-16y+20